OpenStudy (anonymous):

Let sin theta = x and pi/2 < theta < pi, what is cos theta and cos 2 theta in terms of x?

OpenStudy (ikram002p):

\[\cos \theta=\sqrt{1-x}\] \[\cos 2\theta = 1-x^2\]

OpenStudy (anonymous):

Very good and thanks so much that's what I kept getting and would proof correctly but the "solve for x" was hard to wrap my thought process around...